OpenStudy (anonymous):

4sin(x)cos(x)+2sin(x)-2cos(x)-1=0 Having trouble solving this one. I was able to solve it partially by adding the 2cos(x)-1 to the other side, then factoring the rest. To get this: 2sin(x)(2cos(x)+1)=2cos(x)+1 then divide both sides by 2cos(x)+1, we get 2sin(x)=1 solve for it solutions between 0

OpenStudy (asnaseer):

instead of dividing both sides by 2cos(x)+1, move that expression from the right-hand-side to the left-hand-side and factorise

OpenStudy (asnaseer):

\[2\sin(x)(2\cos(x)+1)=2\cos(x)+1\]\[\therefore 2\sin(x)(2\cos(x)+1)-(2\cos(x)+1)=0\]

OpenStudy (asnaseer):

factorise that expression

OpenStudy (asnaseer):

by dividing both sides by 2cos(x)+1 you were effectively throwing away a set of solutions

OpenStudy (mathmale):

@asnaseer : good guidance on your part. Please let @physicscrap try the actual factoring and then respond. Thanks.

OpenStudy (asnaseer):

that was my plan @mathmale :)

OpenStudy (asnaseer):

otherwise the asker will never learn :)

OpenStudy (anonymous):

wow that was fast, thanks guys. got it.

OpenStudy (asnaseer):

yw :)